An equilibrium mixture of H2, I2, and HI at 458 ∘C contains 0.112 molH2, 0.112 molI2, and 0.775 molHI in a 5.00-L vessel.

Part A: What are the equilibrium partial pressure of HI when equilibrium is reestablished following the addition of 0.200 mol of HI? Express your answer to four significant figures and include the appropriate units.

Part B:What are the equilibrium partial pressure of I2 when equilibrium is reestablished following the addition of 0.200 mol of HI? Express your answer to three significant figures and include the appropriate units.

Part C:What are the equilibrium partial pressure of H2 when equilibrium is reestablished following the addition of 0.200 mol of HI? Express your answer to three significant figures and include the appropriate units.
Answer #1

Solo Given Temperature T+ 158°C = 15872:13 - 220K volume, VE 5.00€ Now we have to write the seation under In the equilibrium(C+0.20 -x) (0.775) (a+h) (6+1 / (012) ___ 22 0.0114845. number of moles of He is ___ Hz = 0.112+ Oic H185 - 2 - = 0.134425using the ideal gay equation Presse encerted on this nigra - P Oililykes*0-0824*71 1 . Pre Wo atm 1.6latm Therefore, the part